Which of these equations is the exponential form of ln ( 1/e ) = –1?
Answer by Edwin McCravy(20086)

The exponential form of 

ln(A) = B

is

A = eB

So the exoponential form of

ln(1/e) = –1

is found by substituting 1/e for A and -1 for B in

A = eB

which is:

1/e = e-1

Edwin